# From Tokens to Thoughts: How LLMs and Humans Trade Compression for Meaning

### v5

@@ Abstract
Humans organize knowledge into compact categories that balance compression with semantic meaning preservation. Large Language Models (LLMs) demonstrate striking linguistic abilities, yet whether they achieve this same balance remains unclear. We apply the Information Bottleneck principle to quantitatively compare how LLMs and humans navigate this compression-meaning trade-off. Analyzing embeddings from 40+ LLMs against classic human categorization benchmarks (Rosch, 1973a; 1975; McCloskey & Glucksberg, 1978), we uncover three key findings. First, LLMs broadly align with human categories but miss fine-grained semantic distinctions crucial for human understanding. Second, LLMs demonstrate aggressive statistical compression, achieving “optimal” information-theoretic efficiency, while humans prioritize contextual richness and adaptive flexibility. Third, encoder models surprisingly outperform decoder models in human alignment, suggesting that generation and understanding rely on distinct mechanisms in current architectures. In addition, training dynamics analysis reveals that conceptual structure develops in distinct phases: rapid initial formation followed by architectural reorganization, with semantic processing migrating from deeper to mid-network layers as models discover more efficient encoding. These divergent strategies, where LLMs optimize for compression and humans for adaptive utility, reveal fundamental differences between artificial and biological intelligence, guiding development toward more human-aligned AI.
@@ 1 The Enigma of Meaning in Large Language Models
“The categories defined by constructions in human languages may vary from one language to the next, but they are mapped onto a common conceptual space, which represents a common cognitive heritage, indeed the geography of the human mind.” Croft (2001) p. 139
Humans excel at organizing knowledge into concepts which are compact categories that achieve remarkable compression while preserving essential meaning (Murphy, 2004). A single word like “bird” compresses information about thousands of species, yet maintains critical semantic properties (can fly, has feathers, lays eggs). This hierarchical organization (robin $\rightarrow$ bird $\rightarrow$ animal; Rosch et al. 1976) represents a fundamental cognitive achievement: balancing efficiency with semantic fidelity.
Large Language Models (LLMs) demonstrate striking linguistic capabilities that suggest semantic understanding (Singh et al., 2024; Li et al., 2024). Yet, a critical question remains unanswered: Do LLMs navigate the compression-meaning trade-off similarly to humans, or do they employ fundamentally different representational strategies? This question matters because true understanding, which goes beyond surface-level mimicry, requires representations that balance statistical efficiency with semantic richness (Tversky, 1977; Rosch, 1973b).
To address this question, we apply Rate-Distortion Theory (Shannon, 1948) and Information Bottleneck principles (Tishby et al., 2000) to systematically compare LLM and human conceptual structures. We digitize and release seminal cognitive psychology datasets (Rosch, 1973b; 1975; McCloskey & Glucksberg, 1978), which are foundational studies that shaped our understanding of human categorization but were previously unavailable in a machine-readable form. These benchmarks, comprising 1,049 items across 34 categories with both membership and typicality ratings, offer unprecedented empirical grounding for evaluating whether LLMs truly understand concepts as humans do. It also offers much better quality data than the current crowdsourcing paradigm.
@@ 2 Research Questions and Scope
Prior work has explored LLM conceptual representations through multiple lenses: relational knowledge (Shani et al., 2023; Misra et al., 2021), interpretable concept extraction (Hoang-Xuan et al., 2024; Maeda et al., 2024), sparse activation patterns (Li et al., 2024), and embedding geometry including hierarchical structures (Park et al., 2024). While insightful, these studies often lack deep, quantitative comparison of the compression-meaning trade-off using information theory against rich human cognitive benchmarks.
Separately, cognitive science has applied information theory to human concept learning (Imel & Zaslavsky, 2024; Tucker et al., 2025; Zaslavsky et al., 2018; Sorscher et al., 2022). For example, Zaslavsky et al. (2018) developed an Information Bottleneck framework for color naming efficiency, later extended to animal taxonomies (Zaslavsky et al., 2019). Yet these cognitive studies typically proceed without connecting to modern LLMs, and tend to focus on a specific domain.
@@ 3.1 Human Baselines: Empirical Data from Seminal Cognitive Science
Rosch (1973): This foundational work (Rosch, 1973a) explored semantic categories as part of the research program leading to prototype theory (Rosch, 1973b). The theory posits that categories organize around ”prototypical” members rather than strict, equally shared features. The dataset includes 48 items in eight common semantic categories (e.g., furniture, bird), with prototypicality rankings (e.g., ‘robin’ as typical bird, ‘bat’ as atypical).
Rosch (1975): Building on prototype theory, Rosch (1975) further detailed how semantic categories are cognitively represented. This work provides typicality ratings for a larger set of 552 items across ten categories (e.g., ‘orange‘ as a prototypical fruit, ‘squash‘ as less so).
McCloskey & Glucksberg (1978): Investigated the ”fuzzy” boundaries of natural categories, showing membership is graded rather than absolute (McCloskey & Glucksberg, 1978). Covers 449 items in 18 categories with both typicality scores and membership certainty ratings (e.g., ‘dress’ is typical clothing, ‘bandaid’ less so).
While originating from different researchers, these datasets share rigorous experimental designs and provide data on both category assignments and item typicality. We aggregated data from these studies, creating a unified benchmark of 1,049 items across 34 categories. This data, which we have digitized and made publicly available (Appendix B.1), offers a high-quality empirical foundation for evaluating the human-likeness of LLMs. [^1] Appendix B.6 shows that polysemy is rare in our data and cannot account for our findings.
@@ 3.2 Large Language Models Under Study
We analyze 40+ diverse LLMs spanning multiple architectures and scales (300M to 72B parameters) to understand how conceptual representation varies across model design choices.
Model Selection. Our study encompasses three architectural paradigms. Encoder models include the BERT family (Devlin et al., 2019; He et al., 2020; Zhuang et al., 2021) and CLIP ViT text encoders (Radford et al., 2021). Decoder models form the majority of our analysis: the Llama family (1B-70B; Touvron et al., 2023a; b; Grattafiori et al., 2024), Gemma variants (2B-27B; Team et al., 2024; 2025), Qwen models (0.5B-72B; Bai et al., 2023; Yang et al., 2024), Phi series (Javaheripi et al., 2023; Abdin et al., 2024; Abouelenin et al., 2025), Mistral-7B (Karamcheti et al., 2021), GPT-2 (Radford et al., 2019), and OLMo-7B (Groeneveld et al., 2024). We also include classic embeddings Word2Vec (Mikolov et al., 2013a; b) and GloVe (Pennington et al., 2014) as baselines.
This diverse selection enables us to disentangle effects of architecture (encoder vs. decoder), scale (300M to 72B), and training objectives (understanding vs. generation). Complete model specifications appear in Appendix B.2.
@@
This dual approach allows us to trace how concepts emerge from basic lexical knowledge to contextualized understanding. Critically, our results prove robust to prompt templates and pooling strategies (Appendix B.3). Moreover, despite substantial vocabulary overlap between model families, neither token count nor tokenization patterns correlate with our results (Appendix B.7).
@@ 4 A Framework for Comparing Compression and Meaning
To quantitatively compare how LLMs and humans navigate the fundamental tension between compact representation and semantic richness, we develop a framework that captures both aspects of conceptual organization. Our approach draws on Rate-Distortion Theory (Shannon, 1948) and the Information Bottleneck principle (Tishby et al., 2000), adapting them to measure the quality of conceptual systems.
@@ 4.1 Theoretical Foundations
Human concepts achieve remarkable efficiency: the word ”bird” compresses knowledge about thousands of species into a single category, yet preserves critical semantic information (can fly, has feathers, lays eggs). This reflects a fundamental trade-off that any conceptual system must navigate:
@@
Rate-Distortion Theory (RDT; Shannon, 1948) formalizes this trade-off for lossy compression. Given data $X$ and compressed representation $\hat{X}$, RDT seeks encodings that minimize:
@@
Information Bottleneck (IB; Tishby et al., 2000) extends this by compressing $X$ into $Z$ while preserving information about relevant variable $Y$:
@@
Our adaptation: For conceptual representation, we lack an external relevance variable $Y$. Instead,“relevance” becomes internal semantic coherence, which means how well categories preserve within-group similarity. We thus combine RDT’s geometric distortion with IB’s information-theoretic compression, yielding our framework where clustering $C$ represents items $X$ by minimizing both the information needed to specify items (compression) and the semantic spread within clusters (distortion).
@@ 4.3 Connecting Framework to Research Questions
[RQ2] Internal Semantic Structure: Do LLMs capture human-like typicality gradients? The Distortion term measures how well clusters preserve semantic coherence. In our typicality analysis, we check whether typical items cluster tightly near centroids while atypical items lie farther away. Low distortion with clear center-periphery structure indicates prototype organization that mirrors human cognitive structure.
@@ 5 An Empirical Investigation of Representational Strategies
Approach: We tested whether LLMs naturally organize our 1,049 items into categories resembling human conceptual structure. Token embeddings were extracted at two levels: (i) static embeddings from input layers (E matrix), representing context-free lexical knowledge; (ii) contextual embeddings from all hidden layers, measured layer-wise to identify peak conceptual alignment. These were clustered using k-means ($K$ matching human category counts) and evaluated against human categories using Adjusted Mutual Information (AMI), Normalized Mutual Information (NMI), and Adjusted Rand Index (ARI) metrics.
Broad Categorical Agreement: All 40+ models achieve significant above-chance alignment (Figure 1 (Left)). Even at baseline, static embeddings show substantial alignment (mean $AMI\approx 0.45$), which contextual processing enhances to peak $AMI\approx 0.55$. This confirms that LLMs encode human-like categorical boundaries. Full NMI and ARI results in Appendices B.8B.9.
@@
Figure 1: LLMs capture categorical boundaries but miss semantic gradients. Left: All 40+ models achieve above-chance AMI with human categories, with encoder architectures (squares) matching or exceeding decoder models 100× larger. Results show the layer with peak AMI score per model, see static and mean scores in Figure A.7. Right: Despite categorical success, models show weak correlations ($\rho<0.2$ for most) with human typicality judgments, revealing divergent representational strategies. This divergence between capturing boundaries (compression) while missing internal structure (meaning) reveals how LLMs and humans fundamentally differ in their representational strategies. We note that encoder models align more than decoder models, but these correlations are still modest. Computed using the static embeddings, see full results in Tables 3-4 in Appendix B.10.
@@
Approach: We test whether LLMs encode human-like typicality gradients, i.e., whether robins are more “birdy” than penguins. For each item, we compute the cosine similarity with its category name using embeddings (e.g., ‘robin’ → ‘bird’). We then compare these similarities with human typicality ratings using Spearman’s correlation coefficient $\rho$ (Wissler, 1905).
We employed two analysis approaches: (i) static-layer analysis using embeddings directly from the input layer; (ii) peak AMI layer analysis using contextual embeddings from the layer that maximized AMI in RQ1. For the peak AMI approach, we extracted category embeddings by replacing items with category names in the same prompt template, ensuring consistent contextualization (see Appendix B.4 for template and pooling robustness analysis).
Weak Typicality Alignment: Correlations between LLM internal organization of concepts and human typicality are modest at best (Figure 1 (Right); Tables 3-4 in Appendix B.10). Static embeddings show weak correlations: BERT achieves $\rho=0.38$ ($p<0.05$), while most decoder models fall below $\rho=0.15$. Even when statistically significant, these correlations indicate limited correspondence with human judgments. This shows that the internal concept geometries of models differ from those of humans, with representation-focused models aligning more closely.
@@
This divergence likely stems from training objectives: models explicitly trained for representation learning appear more effective at capturing semantic category relationships in their embeddings, while modern decoder-only models, which optimized primarily for next-token prediction, show consistently lower static-layer correlations. The pattern holds across model scales, suggesting architectural design matters more than size for capturing semantic gradients.
@@
Results: Our analysis reveals three key patterns (Figure 2; full results in Appendix B.13):
LLMs Achieve Superior Statistical Efficiency: Both measures reveal stark differences between LLMs and humans (Figure 2; full results in Appendix B.13):
Higher human entropy. Human concepts consistently exhibit higher mean cluster entropy than LLM-derived clusters at comparable $K$ values, indicating less statistical compactness but greater internal diversity.
@@
Architectural differences. The Complexity-Distortion plot (Figure 2a) reveals that encoder models (BERT, ViT, classic models) consistently achieve superior trade-offs. Their distortion at any given complexity is lower compared to decoder models. This advantage persists across both static and contextual embeddings.
Statistical Optimality Versus Cognitive Utility: This divergence reveals fundamental differences in optimization pressures. LLMs, trained on massive text corpora, develop maximally efficient statistical representations that minimize redundancy and internal variance. Human conceptual systems, though appearing “suboptimal” by information-theoretic measures, serve distinct cognitive needs: flexible generalization, causal reasoning, and robust communication under neural constraints.
Our analysis reveals that compression efficiency does not predict functional capability. We find no correlation between $\mathcal{L}$ scores and downstream performance (r=0.20r=-0.20, $\rho=0.51$ on MMLU; Appendix B.14). This suggests that apparent human “inefficiency” reflects optimization for cognitive flexibility rather than statistical compression. While LLMs excel at compact representation, they may sacrifice the semantic richness essential for human-like understanding.
The consistent architectural patterns we observe raise fundamental questions: Do understanding and generation require distinct computational strategies? The superiority of representation-focused models suggests that current unified architectures may be conflating two fundamentally different cognitive tasks.
(a) Complexity-distortion trade-off across architectures
(b) $\mathcal{L}$ objective comparison with human categories
Figure 2: Divergent optimization strategies: LLMs achieve superior information-theoretic efficiency while humans preserve semantic richness. (a) Encoder models (BERT, ViT, classic embeddings) consistently achieve lower distortion than decoder models at any given complexity level. (b) All LLM-derived clusters achieve lower $\mathcal{L}$ values than human categories (dashed line), indicating more “optimal” compression-distortion balance. Data from Rosch (1975).

### v6

@@ Abstract
Humans organize knowledge into compact conceptual categories that balance compression with semantic richness. Large Language Models (LLMs) exhibit impressive linguistic abilities, but whether they navigate this same compression-meaning trade-off remains unclear. We apply an Information Bottleneck framework to compare human conceptual structure with embeddings from 40+ LLMs using classic categorization benchmarks(rosch1973internal; rosch1975cognitive; mccloskey1978natural). We find that LLMs broadly align with human category boundaries, yet fall short on fine-grained semantic distinctions. Unlike humans, who maintain “inefficient” representations that preserve contextual nuance, LLMs aggressively compress, achieving more optimal information-theoretic compression at the cost of semantic richness. Surprisingly, encoder models outperform much larger decoder models in human alignment, suggesting that understanding and generation rely on distinct representational mechanisms. Training-dynamics analysis reveals a two-phase trajectory: rapid initial concept formation followed by architectural reorganization, during which semantic processing migrates from deep to mid-network layers as the model discovers increasingly efficient, sparser encodings. These divergent strategies, where LLMs optimize for compression and humans for adaptive utility, reveal fundamental differences between artificial and natural intelligence. This highlights the need for models that preserve the conceptual “inefficiencies” essential for human-like understanding.
@@ 1 The Enigma of Meaning in Large Language Models
“The categories defined by constructions in human languages may vary from one language to the next, but they are mapped onto a common conceptual space, which represents a common cognitive heritage, indeed the geography of the human mind.” croft2001radical p. 139
Humans excel at organizing knowledge into concepts which are compact categories that achieve remarkable compression while preserving essential meaning (murphy2004big). A single word like “bird” compresses information about thousands of species, yet maintains critical semantic properties (can fly, has feathers, lays eggs). This hierarchical organization (robin $\rightarrow$ bird $\rightarrow$ animal; rosch1976structural) represents a fundamental cognitive achievement: balancing efficiency with semantic fidelity.
Large Language Models (LLMs) demonstrate striking linguistic capabilities that suggest semantic understanding (singh2024rethinking; li2024geometry). Yet, a critical question remains unanswered: Do LLMs navigate the compression-meaning trade-off similarly to humans, or do they employ fundamentally different representational strategies? This question matters because true understanding, which goes beyond surface-level mimicry, requires representations that balance statistical efficiency with semantic richness (tversky1977features; rosch1973prototype).
To address this question, we apply Rate-Distortion Theory (shannon1948mathematical) and Information Bottleneck principles (tishby2000information) to systematically compare LLM and human conceptual structures. We digitize and release seminal cognitive psychology datasets (rosch1973prototype; rosch1975cognitive; mccloskey1978natural), which are foundational studies that shaped our understanding of human categorization but were previously unavailable in a machine-readable form. These benchmarks, comprising 1,049 items across 34 categories with both membership and typicality ratings, offer unprecedented empirical grounding for evaluating whether LLMs truly understand concepts as humans do. It also offers much better quality data than the current crowdsourcing paradigm.
@@ 2 Research Questions and Scope
Prior work has explored LLM conceptual representations through multiple lenses: relational knowledge (shani2023towards; misra2021language), interpretable concept extraction (hoang2024llm; maeda2024decomposing), sparse activation patterns (li2024geometry), and embedding geometry including hierarchical structures (park2024geometry). While insightful, these studies often lack deep, quantitative comparison of the compression-meaning trade-off using information theory against rich human cognitive benchmarks.
Separately, cognitive science has applied information theory to human concept learning (imel2024optimal; tucker2025towards; zaslavsky2018efficient; sorscher2022neural). For example, zaslavsky2018efficient developed an Information Bottleneck framework for color naming efficiency, later extended to animal taxonomies (zaslavsky2019semantic). Yet these cognitive studies typically proceed without connecting to modern LLMs, and tend to focus on a specific domain. One notable example is wu2025building, which examined abstraction transfer in humans and LLMs using a behavioral and cognitive modeling level. Our work is different in the sense that it analyzes how information is preserved or distorted inside LLM embedding spaces under controlled clustering transformations.
@@
Figure 1 overviews the data generation and analyses. Human data was collected by asking whether an item i (e.g., chair) is a good example of the category C (furniture). These ratings are aggregated into ranked similarity profiles for each category. Models generate analogous scores using their embeddings. We then compute three metrics: [RQ1] Mutual Information to assess category recoverability, [RQ2] Spearman correlation to measure alignment with human typicality structure, and [RQ3] a rate-distortion objective capturing the trade-off between representation complexity and meaning preservation.
Figure 1: Overview of the data generation and analyses. Human data was collected by asking whether an item i (e.g., chair) is a good example of the concept C (furniture). These ratings are aggregated into ranked similarity profiles for each category. Models generate analogous scores using their embeddings. We then compute three metrics: [RQ1] MI to assess category recoverability, [RQ2] Spearman correlation to measure alignment with human internal typicality structure, and [RQ3] a rate-distortion objective capturing the trade-off between representation complexity and meaning preservation.
@@ 3.1 Human Baselines: Empirical Data from Seminal Cognitive Science
Rosch (1973): This foundational work (rosch1973internal) explored semantic categories as part of the research program leading to prototype theory (rosch1973prototype) [^1] Prototype theory is only one account of how humans form concepts; exemplar theory offers an alternative based on stored instances. We do not aim to adjudicate between theories, but use this framework because it provides structured data suitable for modeling. Our computational analysis remains compatible with alternative accounts.. The theory posits that categories organize around “prototypical” members rather than strict, equally shared features. The dataset includes 48 items in eight common semantic categories (e.g., furniture, bird), with prototypicality rankings (e.g., ‘robin’ as typical bird, ‘bat’ as atypical).
Rosch (1975): Building on prototype theory, rosch1975cognitive further detailed how semantic categories are cognitively represented. This work provides typicality ratings for a larger set of 552 items across ten categories (e.g., ‘orange‘ as a prototypical fruit, ‘squash‘ as less so).
McCloskey & Glucksberg (1978): Investigated the ”fuzzy” boundaries of natural categories, showing membership is graded rather than absolute (mccloskey1978natural). Covers 449 items in 18 categories with both typicality scores and membership certainty ratings (e.g., ‘dress’ is typical clothing, ‘bandaid’ less so).
While originating from different researchers, these datasets share rigorous experimental designs and provide data on both category assignments and item typicality. We aggregated data from these studies, creating a unified benchmark of 1,049 items across 34 categories. This data, which we have digitized and made publicly available (Appendix C), offers a high-quality empirical foundation for evaluating the human-likeness of LLMs. [^2] Appendix I shows that polysemy is rare in our data and cannot account for our findings.
@@ 3.2 Large Language Models Under Study
We analyze 40+ diverse LLMs spanning multiple architectures and scales (300M to 72B parameters) to understand how conceptual representation varies across model design choices. We note that our analysis requires access to LLMs’ embeddings, rather than just output. Thus, we are unable to use any closed-source frontier models such as GPT-5 and Claude.
Model Selection. Our study encompasses three architectural paradigms. Encoder models include the BERT family (devlin-etal-2019-bert; he2020deberta; zhuang-etal-2021-robustly) and CLIP ViT text encoders (radford2021learning). Decoder models form the majority of our analysis: the Llama family (1B-70B; touvron2023llama; touvron2023llama2; grattafiori2024llama), Gemma variants (2B-27B; team2024gemma; team2025gemma), Qwen models (0.5B-72B; bai2023qwen; yang2024qwen2), Phi series (javaheripi2023phi; abdin2024phi; abouelenin2025phi), Mistral-7B (karamcheti2021mistral), GPT-2 (radford2019language), and OLMo-7B (olmo2024). We also include classic embeddings Word2Vec (mikolov-etal-2013-efficient; mikolov-etal-2013-distributed) and GloVe (pennington-etal-2014-glove) as baselines.
This diverse selection enables us to disentangle effects of architecture (encoder vs. decoder), scale (300M to 72B), and training objectives (understanding vs. generation). Note that encoder-only models are less represented (and are smaller) since recent LLM development has prioritized decoder-only architectures. Complete model specifications appear in Appendix D.
@@
This dual approach allows us to trace how concepts emerge from basic lexical knowledge to contextualized understanding. Critically, our results prove robust to prompt templates and pooling strategies (Appendix E). Moreover, despite substantial vocabulary overlap between model families, neither token count nor tokenization patterns correlate with our results (Appendix J).
@@ 4 A Framework for Comparing Compression and Meaning
To quantitatively compare how LLMs and humans navigate the fundamental tension between compact representation and semantic richness, we develop a framework that captures both aspects of conceptual organization. Our approach adapts Rate-Distortion Theory (shannon1948mathematical) and the Information Bottleneck principle (tishby2000information) to measure the quality of conceptual systems.
@@ 4.1 Theoretical Foundations
Human concepts achieve remarkable efficiency: the word “bird” compresses knowledge about thousands of species into a single category, yet preserves critical semantic information (can fly, has feathers, lays eggs). This reflects a fundamental trade-off that any conceptual system must navigate:
@@
Rate-Distortion Theory (RDT; shannon1948mathematical) formalizes this trade-off for lossy compression. Given data $X$ and compressed representation $\hat{X}$, RDT seeks encodings that minimize:
@@
Information Bottleneck (IB; tishby2000information) extends this by compressing $X$ into $Z$ while preserving information about relevant variable $Y$:
@@
Our adaptation: For conceptual representation, we lack an external relevance variable $Y$. Instead,”relevance” becomes internal semantic coherence, which means how well categories preserve within-group similarity. We thus combine RDT’s geometric distortion with IB’s information-theoretic compression, yielding our framework where clustering $C$ represents items $X$ by minimizing both the information needed to specify items (compression) and the semantic spread within clusters (distortion).
@@ 4.3 Connecting Framework to Research Questions
[RQ2] Internal Semantic Structure: Do LLMs capture human-like typicality signal? The Distortion term measures how well clusters preserve semantic coherence. In our typicality analysis, we check whether typical items cluster tightly near centroids while atypical items lie farther away. Low distortion with clear center-periphery structure indicates prototype organization that mirrors human cognitive structure.
@@ 5 An Empirical Investigation of Representational Strategies
Approach: We tested whether LLMs naturally organize our 1,049 items into categories resembling human conceptual structure. Token embeddings were extracted at two levels: (i) static embeddings from input layers (E matrix), representing context-free lexical knowledge; (ii) contextual embeddings from all hidden layers, measured layer-wise to identify peak conceptual alignment. These were clustered using k-means ($K$ matching human category counts) and evaluated against human categories using Adjusted Mutual Information (AMI), Normalized Mutual Information (NMI), and Adjusted Rand Index (ARI) metrics. NMI quantifies how much information is shared between the model-derived clusters and the human-labeled categories; AMI refines this measure by correcting for the amount of overlap that would be expected by chance; and ARI assesses the degree of agreement between the two partitions while explicitly accounting for random assignments, providing a complementary view of clustering accuracy.
Broad Categorical Agreement: All 40+ models achieve significant above-chance alignment (Figure 2 (Left)). Even at baseline, static embeddings show substantial alignment (mean $AMI\approx 0.45$), which contextual processing enhances to peak $AMI\approx 0.55$. This confirms that LLMs encode human-like categorical boundaries. Full NMI and ARI results in Appendices K-M.
@@
Figure 2: LLMs capture categorical boundaries (RQ1 AMI scores) but miss internal geometry (RQ2 Spearman correlations). Left: All 40+ models achieve above-chance AMI with human categories, with encoder architectures (squares, circles, and Xs) matching or exceeding decoder models 100× larger (stars). Results show the layer with peak AMI score per model (see static and mean scores in Figure A.8). Right: Despite categorical success, models show weak correlations ($\rho<0.2$ for most) with human typicality judgments, revealing divergent representational strategies. This divergence between capturing boundaries (compression) while missing internal structure (meaning) reveals how LLMs and humans fundamentally differ in their representational strategies. We note that encoder models align more than decoder models, but these correlations are still modest. Computed using the static embeddings, see full results in Tables 3-4 in Appendix N.
@@
Approach: We test whether LLMs encode human-like typicality signals, i.e., whether robins are more “birdy” than penguins. For each item, we compute the cosine similarity with its category name using embeddings (e.g., ‘robin’ → ‘bird’; $cosine_{sim}(E(robin),E(bird)$). We then compare these similarities with human typicality ratings using Spearman’s correlation coefficient $\rho$ (wissler1905spearman).
We employed two analysis approaches: (i) static-layer analysis using embeddings directly from the input layer; (ii) peak AMI layer analysis using contextual embeddings from the layer that maximized AMI in RQ1. For the peak AMI approach, we extracted category embeddings by replacing items with category names in the same prompt template, ensuring consistent contextualization (see Appendix F for template and pooling robustness analysis).
Weak Typicality Alignment: Correlations between LLM internal organization of concepts and human typicality are modest at best (Figure 2 (Right); Tables 3-4 in Appendix N). Static embeddings show weak correlations: BERT achieves $\rho=0.38$ ($p<0.05$), while most decoder models fall below $\rho=0.15$. Even when statistically significant, these correlations indicate limited correspondence with human judgments. This shows that the internal concept geometries of models differ from those of humans, with representation-focused models aligning more closely.
@@
This divergence likely stems from training objectives: models explicitly trained for representation learning appear more effective at capturing semantic category relationships in their embeddings, while modern decoder-only models, which optimized primarily for next-token prediction, show consistently lower static-layer correlations. The pattern holds across model scales, suggesting architectural design matters more than size for capturing fine-grained semantic similarity.
@@
Results: Our analysis reveals three key patterns (Figure 3; full results in Appendix Q):
LLMs Achieve Superior Statistical Efficiency: Both measures reveal stark differences between LLMs and humans (Figure 3; full results in Appendix Q):
Higher human entropy. Human concepts consistently exhibit higher cluster entropy than LLM clusters at comparable $K$ values, indicating less statistical compactness but greater internal diversity.
@@
Architectural differences. The Complexity-Distortion plot (Figure 3a) reveals that encoder models (BERT, ViT, classic models) achieve superior trade-offs. Their distortion at any given complexity is lower compared to decoder models, across both static and contextual embeddings.
Statistical Optimality Versus Cognitive Utility: This divergence reveals fundamental differences in optimization pressures. LLMs, trained on massive text corpora, develop maximally efficient statistical representations that minimize redundancy and internal variance. Although human conceptual systems may look suboptimal under information-theoretic measures, prior work indicates that they are structured to support goals such as flexible generalization and causal reasoning rather than maximal compression (murphy2004big). Thus, the differing pressures shaping human and LLM representations help explain this apparent suboptimality.
Our analysis reveals that compression efficiency does not predict functional capability. We find no correlation between $\mathcal{L}$ scores and downstream performance (r=0.20r=-0.20, $\rho=0.51$ on MMLU; Appendix S). This suggests that apparent human “inefficiency” reflects optimization for cognitive flexibility rather than statistical compression. While LLMs excel at compact representation, they may sacrifice the semantic richness essential for human-like understanding.
The consistent architectural patterns observed raise fundamental questions: Do understanding and generation require distinct computational strategies? The superiority of representation-focused models suggests that current architectures may be conflating two fundamentally different cognitive tasks.
(a) Complexity-distortion trade-offs.
(b) $\mathcal{L}$ objective comparison with human categories.
Figure 3: Divergent optimization strategies: LLMs achieve superior information-theoretic efficiency while humans preserve semantic richness. (a) Encoder models (BERT, ViT, classic embeddings) consistently achieve lower distortion than decoder models at any given complexity level. (b) All LLM-derived clusters achieve lower $\mathcal{L}$ values than human categories (dashed line), indicating more “optimal” compression-distortion balance. Data from rosch1975cognitive.